OpenStudy (anonymous):

How would I solve for v -3/v+4=7

OpenStudy (anonymous):

\[\frac{-3}{v+4}=7\]?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

multiply by \(v+4\) to get rid of the fraction don't forget parentheses \[-3=7(v+4)\]

OpenStudy (anonymous):

multiply 7 by v+4 and you get -3 =7v + 28

OpenStudy (anonymous):

subtract 28 from both sides, divide by 7 thus getting -31/7 = v